OpenStudy (anonymous):

Solve the equation for y. 4x – 5y = 9

OpenStudy (gorv):

add 5 y both side

OpenStudy (gorv):

4x-5y+5y=9+5y 4x=9+5y

OpenStudy (gorv):

subtract 9 from both side 4x-9=5y+9-9 4x-9=5y

OpenStudy (gorv):

divide both side by 5 (4x-9)/5=y

OpenStudy (zpupster):

subtract 4x both sides and divide by -5 remember ti divide each tem by -5

OpenStudy (anonymous):